OpenStudy (anonymous):

Evaluate 8^2/a^3 for a = 2. 2 4 8 16

Parth (parthkohli):

What is \(8^2\)?

OpenStudy (anonymous):

substutute for a

Parth (parthkohli):

@Akeller27 Are you there?

Parth (parthkohli):

Do you know what \(8^2\) is?

OpenStudy (anonymous):

yes im here and no

Parth (parthkohli):

Okay, what is \(8 \times 8\)?

OpenStudy (anonymous):

8^2 = 8 * 8 = 64

OpenStudy (anonymous):

2^3 = 2* 2 * 2 = 8

OpenStudy (anonymous):